1.) a.) Suppose the dominant allele frequency is 0.15, what is the number of homozygous dominant...

1.)

a.) Suppose the dominant allele frequency is 0.15, what is the number of homozygous dominant individuals you would expect to find in a population of 100 individuals?

b.) Suppose you are trying to determine the allele frequency of hair texture in a group of 50 individuals. You notice 15 individuals have curly hair (HH), 25 have wavy hair (Hh), and 10 have straight hair (hh). What is the allele frequency of the dominant allele, "H"?
Answer #1

1. a) According to Handy- weinberg low, P + 2pq+q² = 1 and p+q=1 Ptq=1 or q = 0,85 T P = 0.15] TP = frequency of dominant alt(HH) Curly Hair = 15 (Hh) Wavy hair: 25 (hh) Straight hair - 10 50 we know that, allele frequency of t = 2X(HH) + (Hh) 2x [